An 5.0 m x25.0m x7.0m. The density of air is 1.20 kg/m
Answer #1

Solution 9)

A) Volume= 35 m * 25 m* 7m = 6125 m^3 = 216302.53 ft^3

=====

B) mass= weight= density* volume= 6125*1.20 =7350 kg = 16217.18 pounds

=======

Acc to HOMEWORKLIB RULES, I can answer only one at a time. Good luck!:)